HETAV PATEL

Grade 11,

tan x tan 2x+ tan 2x tan 3x+........................+tan nx tan(n+1)x

tan x tan 2x+ tan 2x tan 3x+........................+tan nx tan(n+1)x

Grade:11

2 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear Divya ,

use the formulae  : tan(a-b) = tan a -tanb /1+tana tanb , we find , tan a tan b  = tana -tan b -tan(a-b) /tan (a-b)

apply to each term we get , expression

 = (tan2x -tanx -tanx + tan3x- tan2x -tan x+ tan4x -tan3x -tanx ......tan(n+1)x -tannx -tanx )/tanx

 = tan(n+1)x -(n+1)tan x /tanx

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Win exciting gifts by answering the questions on Discussion Forum..

Harsha Kumari
5 years ago
tanx + tan2x+ tan3x=0
or, tanx+ tan2x= - tan3x
or, (sinx/cosx)+(sin 2x /Cos 2x)=-(sin 3x /cos 3x)
or, (sin x *cos 2 X + cos x *sin 2x)/cos x * cos 2x = -sin 3x / cos 3x
or, sin( 2x+x)* cos 3x= - cos x *cos 2x *sin 3x
or, sin 3x*cos 3x+ cos x * cos 2x *sin 3x =0
or, sin 3x ( cos 3x + cos x *cos 2x)=0
or, sin 3x (cos ( 2x + x)+ cos x * cos 2x)=0
or, sin 3x (cos x *cos 2x - sin x *sin 2x+ cos x *cos 2x) =0
or, -sin 3x *sin x * sin 2x=0
 
Either sin3x=0
i.e, 3x=nπ
i.e, X=nπ/3
 
or, sin 2x=0
i.e, 2x=nπ
i.e, X= nπ/2
 
or, sin x= 0
i.e, X=nπ
Required solution is X= nπ/3